OpenStudy (anonymous):

(2)(^4 sqrt3) * (4)(^4 sqrt 7)

OpenStudy (anonymous):

\[2\sqrt[4]{3} * 4\sqrt[4]{7}\]

OpenStudy (jdoe0001):

same root, thus just combine and multiply the radicands

OpenStudy (anonymous):

\[if we have \sqrt[n]{x} * \sqrt[n]{y} = \sqrt[n]{x * y}\]

OpenStudy (anonymous):

\[\sqrt[4]{21}\]

OpenStudy (anonymous):

Correct !!

OpenStudy (anonymous):

but what about the 2 and 4

OpenStudy (anonymous):

So what would be the final answer ?

OpenStudy (anonymous):

2 * 4 * \[\sqrt[4]{21}\]

OpenStudy (anonymous):

\[8\sqrt[4]{21} ?\]

OpenStudy (anonymous):

Correct !!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

u r most welcome :)